# Charactarisation of distal actions of automorphisms on the space of one-parameter subgroups of Lie groups

### v1

@@ 2. Actions of automorphisms of $G$ on ${\mathrm{Sub}}_{G}$ and ${\mathrm{Sub}}^{p}_{G}$
Let $G$ be a locally compact second countable group. A sequence $\{H_{n}\}\subset{\rm Sub}_{G}$ converges to $H\in{\rm Sub}_{G}$ if and only if the following hold:
@@ 2. ${\rm(II)}$
The following lemma is well-known. We include a short proof for the sake of completeness.
@@
Let $G$ be a second countable Lie group. Suppose $\{H_{n}\}_{n\in{\mathbb{N}}}$ in ${\rm Sub}_{G}$ is such that it converges to $H\in{\rm Sub}_{G}$. If $H$ is discrete, then $H_{n}$ is discrete for all large $n$. In particular, the set of discrete subgroups of $G$ is open in ${\rm Sub}_{G}$.
Proof.
It is enough to show that $H_{n}^{0}$ is trivial for all large $n$. As $G^{0}$ is open in $G$, we get that $H_{n}\cap G^{0}\to H\cap G^{0}$ and the latter is discrete. As $H_{n}^{0}\subset G^{0}$, we may replace $H_{n}$ by $H_{n}\cap G^{0}$, $H$ by $H\cap G^{0}$, and $G$ by $G^{0}$ and assume that $G$ is connected. As $G$ is a Lie group, there exists a neighbourhood $U$ of the identity $e$ in $G$ such that the only subgroup contained in $U$ is the trivial subgroup $\{e\}$. As $H$ is discrete, there exists a neighbourhood $W$ of the identity $e$ in $G$ such that $H\cap W=\{e\}$. Since $G$ is locally compact and Hausdorff, there exist an open set $V$ and a compact set $K$ such that $e\in V\subset K\subset U\cap W$. As $G$ is connected, $K\cap V^{c}$ is a nonempty compact set and
$H\cap(K\cap V^{c})\subset H\cap W\cap V^{c}=\emptyset,$
and hence $H\in U_{1}(K\cap V^{c})$ which is an open set in ${\rm Sub}_{G}$. As $H_{n}\to H$ in ${\rm Sub}_{G}$, we get that for all large $n$, $H_{n}\in U_{1}(K\cap V^{c})$, i.e. $H_{n}\cap(K\cap V^{c})=\emptyset$, and hence $H_{n}\subset V\sqcup K^{c}$. Thus for all large $n$, $H_{n}^{0}\subset V\subset U$, and it follows that $H_{n}^{0}=\{e\}$, and hence $H_{n}$ is discrete. The second assertion easily follows from the first. ∎
Recall that an automorphism of $G$ is a homomorphism and a homeomorphism, and ${\rm Aut}(G)$ denotes the group of all automorphisms of $G$. There is a natural group action of ${\rm Aut}(G)$ on ${\rm Sub}_{G}$ defined as follows:
${\rm Aut}(G)\times{\rm Sub}_{G}\to{\rm Sub}_{G},\ (T,H)\mapsto T(H);\ T\in{\rm Aut}(G),\ H\in{\rm Sub}_{G}.$
It is easy to see that the map $H\mapsto T(H)$ defines a homeomorphism of ${\rm Sub}_{G}$ for every $T\in{\rm Aut}(G)$, and the corresponding map is $\phi:{\rm Aut}(G)\rightarrow\mbox{Homeo}({\rm Sub}_{G})$, where $\phi(T)(H)=T(H)$. It is easy to check that $\phi$ is a group homomorphism.
Recall that for a locally compact Hausdorff group $G$, if ${\rm Aut}(G)$ is equipped with the compact-open topology, then
${\rm Aut}(G)\times G\rightarrow G,(T,x)\mapsto T(x),T\in{\rm Aut}(G),x\in G,$
is continuous. If we consider the modified compact-open topology on ${\rm Aut}(G)$, it is a topological group [42]. In Lemma 2.4 of [40], it is proved that the action of the group ${\rm Aut}(G)$ on ${\rm Sub}_{G}$ is continuous, implying that $\phi$ is continuous. Note that for a connected Lie group $G$, the compact-open topology and the modified compact-open topology on ${\rm Aut}(G)$ are same. In particular, every $T$ which generates a relatively compact group in ${\rm Aut}(G)$ acts distally on ${\rm Sub}_{G}$ (see e.g. [30], Lemma 4.2).
Let $G$ be a connected Lie group with the Lie algebra ${\mathfrak{g}}$. There is an exponential map $\exp:{\mathfrak{g}}\to G$ which is a continuous map and its restriction to a small open neighbourhood of 00 in ${\mathfrak{g}}$ is a homeomorphism onto an open neighbourhood of the identity $e$ in $G$. Given $T\in{\rm Aut}(G),$ there exists a corresponding Lie algebra automorphism ${\rm d\,}T:{\mathfrak{g}}\to{\mathfrak{g}}$ such that $\exp\circ\,{\rm d\,}T=T\circ\,\exp$. The map ${\rm d\,}:{\rm Aut}(G)\rightarrow{\rm Aut}({\mathfrak{g}})$, $T\mapsto{\rm d\,}T$ is an injective continuous homomorphism and its image is closed, and in fact, ${\rm d\,}$ is a homeomorphism onto its image (cf. [9], see also [21]) It is known that ${\rm Aut}({\mathfrak{g}})$ is a closed subgroup of ${\rm GL}({\mathfrak{g}})$, and we can view ${\rm Aut}(G)$ as a closed subgroup of ${\rm GL}({\mathfrak{g}})$. The topology on ${\rm Aut}(G)$ when considered as a subspace topology of ${\rm GL}({\mathfrak{g}})$ coincides with the (modified) compact-open topology. We refer the reader to [22] for the basic structure theory of Lie groups.
Let $G$ be a connected Lie group. A (continuous real) one-parameter subgroup in $G$ is the image of a continuous homomorphism from $\mathbb{R}$ to $G$. Any one-parameter subgroup is of the form $\{\exp tX\}_{t\in{\mathbb{R}}}$ for some $X$ in the Lie algebra ${\mathfrak{g}}$ of $G$. Moreover, if $H$ is a one-parameter subgroup in $G$, then so is $T(H)$ for any $T\in{\rm Aut}(G)$. Let
@@
Then ${\mathcal{P}}_{1}(G)\subset{\rm Sub}_{G}$ and it is invariant under the action of ${\rm Aut}(G)$. Let ${\rm Sub}^{p}_{G}$ denote the closure (in ${\rm Sub}_{G}$) of the set ${\mathcal{P}}_{1}(G)$ of all closed one-parameter subgroups in $G$, i.e.
${\rm Sub}^{p}_{G}:=\overline{{\mathcal{P}}_{1}(G)}.$
Then ${\rm Sub}^{p}_{G}$ is compact and it is also invariant under the action of ${\rm Aut}(G)$.
Note that if any one-parameter subgroup in a locally compact group is not closed, then its closure is compact and isomorphic to an $n$ dimensional torus $\mathbb{T}^{n}=(\mathbb{S}^{1})^{n}$ for some $n\in{\mathbb{N}}$ with $n\geq 2$. If it is closed and bounded (compact), then it is isomorphic to $\mathbb{S}^{1}$. If it is unbounded (not relatively compact), then it is closed and isomorphic to ${\mathbb{R}}$.
For brevity, for a set $\{x_{t}\}_{t\in{\mathbb{R}}}$ which may be a continuous one-parameter set or a one-parameter subgroup, we may just write $\{x_{t}\}$, i.e. we may omit the parameter set if there is no ambiguity.
Now we prove some elementary results about the structure of ${\rm Sub}^{p}_{G}$. The following lemma shows that if $G$ contains any non-trivial compact connected abelian subgroup and the dimension of $G$ is at least 2, then ${\mathcal{P}}_{1}(G)$ is not closed in ${\rm Sub}_{G}$.
@@
Let $G$ be a connected Lie group and let ${\rm Sub}^{p}_{G}$ be the closure in ${\rm Sub}_{G}$ of all closed one-parameter subgroups. Then the following holds:
@@
Let $K$ be a compact connected abelian subgroup. Then the set $F$ of all finite order elements in $K$ is dense in $K$ and each such element of $F$ is contained in a closed one-parameter subgroup of $K$. Since $K$ is monothetic (i.e. it has a dense cyclic subgroup), we get that $K$ is a limit of a sequence of compact one-parameter subgroups of $K$ in ${\rm Sub}_{K}$. Therefore, $K\in{\rm Sub}^{p}_{G}$. Thus (1) holds.
Let $\{x_{t}\}\in{\mathcal{P}}_{1}(G)\subset{\rm Sub}^{p}_{G}$ be such that each $x_{t}$ centralises $K$. If $\{x_{t}\}$ is compact, then $\{x_{t}\}K$ is a compact connected abelian subgroup of $G$ and hence $\{x_{t}\}K\in{\rm Sub}^{p}_{G}$ from (1). Now suppose $\{x_{t}\}$ is isomorphic to ${\mathbb{R}}$. Let $k\in K$ be such that it generates a dense cyclic subgroup in $K$. Since $K$ is exponential, there exists a one-parameter subgroup $\{k_{t}\}$, which is dense in $K$ and $k=k_{1}$. Let $\{n_{m}\}$ be such that $k^{n_{m}}\to k$. Let $A_{m}=\{a_{t}(m)=x_{t/n_{m}^{2}}k_{t}\}_{t\in{\mathbb{R}}}$. Then each $A_{m}$ is a closed one-parameter subgroup in $G$, and passing to a subsequence if necessary, we get that $A_{m}\to A$ for some subgroup $A\in{\rm Sub}^{p}_{G}$. Note that $\{x_{t}\}\times K$ is a closed subgroup of $G$ and it contains each $A_{m}$. Therefore, $A\subset\{x_{t}\}\times K$. Since $a_{n_{m}}(m)=x_{1/n_{m}}k_{n_{m}}=x_{1/n_{m}}k^{n_{m}}\to k$, we get that $k\in A$. Therefore, $K\subset A$. For a fixed $t$, let $b_{m}=a_{tn_{m}^{2}}(m)=x_{t}k_{tn_{m}^{2}}\in A_{m}$, $m\in{\mathbb{N}}$. Passing to a subsequence of $\{b_{m}\}$ if necessary, we get that $b_{m}\to x_{t}h$ for some $h\in K$, and hence, $x_{t}\in A$. Now $\{x_{t}\}\subset A$. Hence, $A=\{x_{t}\}\times K$ and (2) holds. ∎
@@
Let $K$ and $\pi$ be as above. Since $K$ is compact and central in $G$, $HK$ is a closed subgroup of $G$ for any closed subgroup $H$ of $G$. Also, if $H_{n}K\to L$ in ${\rm Sub}_{G}$, then since ${\rm Sub}_{G}$ and $K$ are compact, it is easy to see that for any limit point $H$ of $\{H_{n}\}$, $L=HK$. Therefore, $\{HK\mid H\in{\rm Sub}_{G}\}$ is closed. Now if $H\in{\rm Sub}^{p}_{G}$, then there exists $\{A_{n}\}\subset{\mathcal{P}}_{1}(G)$, such that $A_{n}\to H$. By Lemma 2.3 (2), $A_{n}K\in{\rm Sub}^{p}_{G}$. As $A_{n}K\to HK$ and ${\rm Sub}^{p}_{G}$ is closed, we get that $HK\in{\rm Sub}^{p}_{G}$. Now arguing as above, it is easy to show that $\{HK\mid H\in{\rm Sub}^{p}_{G}\}$ is a closed subset of ${\rm Sub}^{p}_{G}$. Moreover, if $T\in{\rm Aut}(G)$ with $T(K)=K$, then $T(HK)=T(H)K$ for any $H\in{\rm Sub}_{G}$. Therefore, the sets mentioned in (1) are invariant under the action of ${\rm Aut}(G)$ if $T(K)=K$. Thus (1) holds.
If $A\in{\mathcal{P}}_{1}(G)$, then $AK$ is closed and $\pi(A)\in{\mathcal{P}}_{1}(G/K)$. Conversely, suppose $B\in{\mathcal{P}}_{1}(G/K$). It is easy to see that $B_{1}:=\pi^{-1}(B)$ is a closed connected nilpotent Lie subgroup of $G$, and any one-parameter subgroup of $B_{1}$, which not contained in $K$, projects onto $B$. Let $U$ be an open neighbourhood of the identity $e$ in $B_{1}$ such that every $x\in U$ is contained in a one-parameter subgroup. As $B_{1}$ is connected, $U\neq K$ and $U\setminus K$ is open. Suppose $B_{1}$ is compact. Being connected and nilpotent, $B_{1}$ is abelian. Since the set of roots of unity in $B_{1}$ is dense and $U\setminus K$ is open in $B_{1}$, we can choose $x\in U\setminus K$ such that $x^{n}=e$ for some $n\in{\mathbb{N}}$. If $B_{1}$ is unbounded, then let $x\in U\setminus K$ be any element. Let $\{x_{t}\}\subset B_{1}$ be a one-parameter subgroup in $B_{1}$ such that $x_{1}=x$. Then $\pi(\{x_{t}\})=\{x_{t}\}K/K=B$. If $B_{1}$ is compact, then $\{x_{t}\}$ is closed as $x_{1}^{n}=e$. If $B_{1}$ is unbounded, then so is $\{x_{t}\}$, and hence it is closed. Therefore, ${\mathcal{P}}_{1}(G/K)=\{\pi(A)\mid A\in{\mathcal{P}}_{1}(G)\}$, i.e. (2) holds.
@@
Let ${\mathcal{S}}(K)=\{H\in{\rm Sub}^{p}_{G}\mid HK=KH=H\}$. It is easy to show that ${\mathcal{S}}(K)$ is closed in ${\rm Sub}^{p}_{G}$, and hence it is compact. Observe that the set ${\mathcal{S}}(K)$ is same as the set $\{HK\mid H\in{\rm Sub}^{p}_{G}\}$. For $H_{1},H_{2}\in{\rm Sub}_{G}$, if $\pi(H_{1})=\pi(H_{2})$, then $H_{1}K=H_{2}K$. Therefore, $H\mapsto\pi(H)$ from ${\mathcal{S}}(K)$ to ${\rm Sub}^{p}_{G/K}$ is injective, and it easily follows from (3) that it is surjective. Also if $H_{n}\to H$ in ${\rm Sub}_{G}$, then $\pi(H_{n})\to\pi(H)$ in ${\rm Sub}_{G/K}$. Since ${\mathcal{S}}(K)$ is compact and ${\rm Sub}^{p}_{G/K}$ is Hausdorff, we have that the map induced by $\pi$ from ${\mathcal{S}}(K)$ to ${\rm Sub}^{p}_{G/K}$ is a homeomorphism. Thus (4) holds. ∎
3. Characterisation of distal actions on the space of one-parameter subgroups
In this section we discuss the action of an automorphism as well as that of a subgroup of automorphisms of a connected Lie group $G$ on ${\rm Sub}^{p}_{G}$ and prove results stated in the introduction. We first set some notations and discuss some preliminary results which are useful.
For convenience, we define a class ${\mathcal{C}^{\prime}}$ of Lie groups as follows:

### v2

@@ 2. Properties and structure of ${\mathrm{Sub}}_{G}$ and ${\mathrm{Sub}}^{p}_{G}$
Let $G$ be a connected Lie group. A sequence $\{H_{n}\}\subset{\rm Sub}_{G}$ converges to $H\in{\rm Sub}_{G}$ if and only if the following hold:
@@ 2. ${\rm(II)}$
The following lemma is well-known (e.g. see Corollary to Lemma 1.2 in [45]; where for a connected Lie group $G$, the conditions for the convergence of a sequence of closed subgroups are equivalent to those in Lemma 2.1 above).
@@
Let $G$ be a connected Lie group. Suppose $\{H_{n}\}_{n\in{\mathbb{N}}}$ in ${\rm Sub}_{G}$ is such that it converges to $H\in{\rm Sub}_{G}$. If $H$ is discrete, then $H_{n}$ is discrete for all large $n$. In particular, the set of discrete subgroups of $G$ is open in ${\rm Sub}_{G}$.
Note that the action of ${\rm Aut}(G)$ on ${\rm Sub}_{G}$ is continuous for such a $G$ (cf. [42], Lemma 2.4). In particular, if $T$ generates a relatively compact group in ${\rm Aut}(G)$, then it acts distally on $G$ as well as on ${\rm Sub}_{G}$ (see e.g. [31], Lemma 4.2 for the latter statement).
A (continuous real) one-parameter subgroup in $G$ is the image of a continuous homomorphism from $\mathbb{R}$ to $G$. Any one-parameter subgroup is of the form $\{\exp tX\}_{t\in{\mathbb{R}}}$ for some $X$ in the Lie algebra ${\mathfrak{g}}$ of $G$. Moreover, if $H$ is a one-parameter subgroup in $G$, then so is $T(H)$ for any $T\in{\rm Aut}(G)$. Let
@@
Now ${\rm Sub}^{p}_{G}$ is the closure (in ${\rm Sub}_{G}$) of the set ${\mathcal{P}}_{1}(G)$, i.e.
${\rm Sub}^{p}_{G}:=\overline{{\mathcal{P}}_{1}(G)},$
which is compact and invariant under the action of ${\rm Aut}(G)$.
Note that if any one-parameter subgroup in a connected Lie group is not closed, then its closure is compact and isomorphic to an $n$ dimensional torus $\mathbb{T}^{n}=(\mathbb{S}^{1})^{n}$ for some $n\in{\mathbb{N}}$ with $n\geq 2$. If it is closed and bounded (compact), then it is isomorphic to $\mathbb{S}^{1}$. If it is unbounded (not relatively compact), then it is closed and isomorphic to ${\mathbb{R}}$.
For $\{x_{t}\}_{t\in{\mathbb{R}}}$, which may be a continuous one-parameter set or a one-parameter subgroup, we may just write $\{x_{t}\}$ for brevity, i.e. we may omit the parameter set if there is no ambiguity.
Now we prove some elementary results about the structure of ${\rm Sub}^{p}_{G}$. The following lemma shows that if $G$ contains any non-trivial compact connected abelian subgroup of dimension at least 2, then ${\mathcal{P}}_{1}(G)$ is not closed in ${\rm Sub}_{G}$.
@@
Let $G$ be a connected Lie group. Then the following hold:
@@
Let $K$ be a compact connected abelian subgroup of $G$. Then the set $F$ of all finite order elements in $K$ is dense in $K$ and each such element of $F$ is contained in a closed one-parameter subgroup of $K$. Since $K$ is monothetic (i.e. it has a dense cyclic subgroup), we get that $K$ is a limit of a sequence of compact one-parameter subgroups of $K$ in ${\rm Sub}_{K}$. Therefore, $K\in{\rm Sub}^{p}_{G}$. Thus (1) holds.
Let $\{x_{t}\}\in{\mathcal{P}}_{1}(G)\subset{\rm Sub}^{p}_{G}$ be such that each $x_{t}$ centralises $K$. If $\{x_{t}\}$ is compact, then $\{x_{t}\}K$ is a compact connected abelian subgroup of $G$ and hence $\{x_{t}\}K\in{\rm Sub}^{p}_{G}$ from (1). Now suppose $\{x_{t}\}$ is isomorphic to ${\mathbb{R}}$. Let $k\in K$ be such that it generates a dense cyclic subgroup in $K$. Since $K$ is exponential, there exists a one-parameter subgroup $\{k_{t}\}$, which is dense in $K$ and $k=k_{1}$. Let $A_{n}=\{a_{t}(n)=x_{t/n}k_{t}\}_{t\in{\mathbb{R}}}$. Then each $A_{n}$ is a closed one-parameter subgroup in $G$, and passing to a subsequence if necessary, we get that $A_{n}\to A$ for some subgroup $A\in{\rm Sub}^{p}_{G}$. Note that $\{x_{t}\}\times K$ is a closed subgroup of $G$ and it contains each $A_{n}$. Therefore, $A\subset\{x_{t}\}\times K$. Let $t\in{\mathbb{R}}$ be fixed. Since $a_{t}(n)=x_{t/n}k_{t}\to k_{t}$ as $n\to\infty$, we get that $k_{t}\in A$, and hence $K\subset A$. Let $b_{n}=a_{tn}(n)=x_{t}k_{tn}\in A_{n}$, $n\in{\mathbb{N}}$. Passing to a further subsequence of $\{b_{n}\}$ if necessary, we get that $b_{n}\to x_{t}h$ for some $h\in K$, and hence, $x_{t}\in A$, as $K\subset A$. Now $\{x_{t}\}\subset A$. Hence, $A=\{x_{t}\}\times K$ and (2) holds. ∎
@@
Let $K$ and $\pi$ be as above. Note that $K\in{\rm Sub}^{p}_{G}$ by Lemma 2.3 If $G=K$, then it is easy to see that $(1-4)$ are trivially satisfied.
Now suppose that $G\neq K$. Since $K$ is compact and central in $G$, $HK$ is a closed subgroup of $G$ for any closed subgroup $H$ of $G$. If $H_{n}K\to L$ in ${\rm Sub}_{G}$, then since both ${\rm Sub}_{G}$ and $K$ are compact, it is easy to show that for any limit point $H$ of $\{H_{n}\}$, $L=HK$. Therefore, $\{HK\mid H\in{\rm Sub}_{G}\}$ is closed. Now if $H\in{\rm Sub}^{p}_{G}$, then there exists $\{A_{n}\}\subset{\mathcal{P}}_{1}(G)$, such that $A_{n}\to H$. By Lemma 2.3 (2), $A_{n}K\in{\rm Sub}^{p}_{G}$. As $A_{n}K\to HK$ and ${\rm Sub}^{p}_{G}$ is closed, we get that $HK\in{\rm Sub}^{p}_{G}$. Now arguing as above, it is easy to show that $\{HK\mid H\in{\rm Sub}^{p}_{G}\}$ is a closed subset of ${\rm Sub}^{p}_{G}$. Moreover, if $T\in{\rm Aut}(G)$ with $T(K)=K$, then $T(HK)=T(H)K$ for any $H\in{\rm Sub}_{G}$. Therefore, the sets mentioned in (1) are invariant under the action of ${\rm Aut}(G)$ if $T(K)=K$. Thus (1) holds.
If $A\in{\mathcal{P}}_{1}(G)$, then $AK$ is closed and $\pi(A)\in{\mathcal{P}}_{1}(G/K)$. Conversely, suppose $B\in{\mathcal{P}}_{1}(G/K$). It is easy to see that $B_{1}:=\pi^{-1}(B)$ is a closed connected nilpotent Lie subgroup of $G$. There exists a neighbourhood $U$ of $e$ such that every element of $U$ is contained in a one-parameter subgroup and $U\setminus K$ is nonempty. Choose any $x\in U\setminus K$, there exists a one-parameter subgroup $\{x_{t}\}$ with $x=x_{1}$. Then $\pi(\{x_{t}\})=B$ and $B_{1}=\{x_{t}\}K$. As $K$ is central in $G$ we get that $B_{1}$ is, in fact, abelian. If $\{x_{t}\}$ is closed, then it belongs to ${\mathcal{P}}_{1}(G)$ and $\pi(\{x_{t}\})=B$. Now suppose it is not closed, then its closure is compact, and hence $B_{1}$ is compact. This implies that the set of torsion elements is dense in $B_{1}$. Since $U\setminus K$ is open, we can choose $y\in U\setminus K$ such that $y^{n}=e$, for some $n\in{\mathbb{N}}\setminus\{1\}$. Let $\{y_{t}\}$ be a one-parameter subgroup in $B_{1}$ such that $y_{1}=y$. Then $\pi(\{y_{t}\})=B$. Here, $\{y_{t}\}$ is closed as $y_{1}^{n}=e$. Therefore, ${\mathcal{P}}_{1}(G/K)=\{\pi(A)\mid A\in{\mathcal{P}}_{1}(G)\}$, i.e. (2) holds.
@@
Let ${\mathcal{S}}(K)=\{H\in{\rm Sub}^{p}_{G}\mid HK=KH=H\}$. It is easy to show that ${\mathcal{S}}(K)$ is closed in ${\rm Sub}^{p}_{G}$, and hence it is compact. Observe that the set ${\mathcal{S}}(K)$ is same as the set $\{HK\mid H\in{\rm Sub}^{p}_{G}\}$. For $H_{1},H_{2}\in{\rm Sub}_{G}$, if $\pi(H_{1})=\pi(H_{2})$, then $H_{1}K=H_{2}K$. Therefore, the map $H\mapsto\pi(H)$ from ${\mathcal{S}}(K)$ to ${\rm Sub}^{p}_{G/K}$ is injective, and it easily follows from (3) that it is surjective. Also if $H_{n}\to H$ in ${\rm Sub}_{G}$, then $\pi(H_{n})\to\pi(H)$ in ${\rm Sub}_{G/K}$. Since ${\mathcal{S}}(K)$ is compact and ${\rm Sub}^{p}_{G/K}$ is Hausdorff, we have that the map induced by $\pi$ from ${\mathcal{S}}(K)$ to ${\rm Sub}^{p}_{G/K}$ is a homeomorphism. Thus (4) holds. ∎
3. Some properties of automorphisms and their actions on $G$ and ${\mathrm{Sub}}^{p}_{G}$
In this section, we discuss some properties and actions of automorphisms of a connected Lie group $G$ on $G$ as well as ${\rm Sub}^{p}_{G}$ and discuss some preliminary results which are useful.
For convenience, we introduce a class ${\mathcal{C}^{\prime}}$ of Lie groups as follows: